Two charges are placed on the x- axis 3uc at x=0 & 5uc at x= 45cm. Where must a third charge q be placed if the force it experience is to be zero?


1
Expert's answer
2021-07-06T13:35:26-0400

"k\frac{q_1q_2}{x^2}-k\frac{q_1q}{x_q^2}=0\to x_q=\sqrt{q\cdot x^2/q_2}=\sqrt{q\cdot 0.45^2/(5\cdot10^{-6})}="


"=201.25\cdot\sqrt{q} \ (m)"